OpenStudy (anonymous):

what is 8x to the power of 3 in a expression

OpenStudy (anonymous):

\[(8x)^3 = 8^3 \times x^3 = 512x^3\]

OpenStudy (anonymous):

it says in simplest form

OpenStudy (anonymous):

@Ankh

OpenStudy (anonymous):

is (8x)^3 the whole expression or part of a bigger expression?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

What is the problem? "Simplify (8x)^3"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then my first reply would be the answer. It's a simple form.

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

How it's simple or how to do it?

OpenStudy (anonymous):

i give up

OpenStudy (anonymous):

You don't know how I got the answer? I can explain it slowly.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Ankh

OpenStudy (anonymous):

Do you know how to do this one? \[(2x)^2\]